# Fairfax, Virginia, USA

### v1

@@ 2.2 Related work
Again, in [1], the authors have shared their ongoing research on the issue of code-mixed social media data and artificial language identification. They discuss a brand-new dataset made up of Facebook comments and posts that exhibit code-mixing between Bengali, English, and Hindi. Additionally, they provide some preliminary research on word-level language recognition using several approaches, such as an unsupervised dictionary-based approach, supervised word-level classification with and without contextual cues, and conditional random fields for sequence labeling.
On the other hand, code mixed data is difficult to find in bulk quantity as it is informal in nature and mostly available in social media posts only. This data shortage has already been addressed by creating synthetic code mixed data to supplement the real code mixed data. Parallel sentences can be used to create code-mixed sentences, as shown by [16] and [2], who also illustrate how utilizing these artificial sentences can increase language model confusion. By employing phrase tables to align and combine components of a parallel sentence, [25] creates CM sentences. [18] also put out a similar approach that creates artificial phrases using parse trees. When optimizing a transformer for a downstream task, [12] provides a novel way to enhance the monolingual source data via multilingual code-switching via random translations.
@@ 4 Results and Analysis
[…]
@@
[…]
@@
[…]

### v2

@@ 2.2 Related work
Again, in [2], the authors have shared their ongoing research on the issue of code-mixed social media data and artificial language identification. They discuss a brand-new dataset made up of Facebook comments and posts that exhibit code-mixing between Bengali, English, and Hindi. Additionally, they provide some preliminary research on word-level language recognition using several approaches, such as an unsupervised dictionary-based approach, supervised word-level classification [17] with and without contextual cues, and conditional random fields for sequence labeling. Similar works include [9], [23] and [11]. Some works also focus on words extracted from image data, which is very likely to be code mixed as well [1]. Such works also need to focus on traditional edge detection methods [25].
On the other hand, code mixed data is difficult to find in bulk quantity as it is informal in nature and mostly available in social media posts only. The only relevant works include SentMix-3L [22], OffMix-3L [10] and TB-OLID [24]. This data shortage has already been addressed by creating synthetic code mixed data to supplement the real code mixed data. Parallel sentences can be used to create code-mixed sentences, as shown by [21] and [3], who also illustrate how utilizing these artificial sentences can increase language model confusion. By employing phrase tables to align and combine components of a parallel sentence, [34] creates CM sentences. [27] also put out a similar approach that creates artificial phrases using parse trees. When optimizing a transformer for a downstream task, [16] provides a novel way to enhance the monolingual source data via multilingual code-switching via random translations.
@@ 4 Results and Analysis
HindiBERTIndicBERTBanglaBERTDistilBERTEmoBERTaroBERTaBERTHingBERTMuRILBanglishBERTmBERTXLM-RTri-Distil-BERTMixed-Distil-BERT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.320.320.320.320.380.380.380.380.390.390.390.390.40.40.40.40.410.410.410.410.410.410.410.410.440.440.440.440.450.450.450.450.450.450.450.450.470.470.470.470.490.490.490.490.510.510.510.510.480.480.480.480.50.50.50.5Weighted Weighted F1 Score
@@
HindiBERTBanglaBERTDistilBERTIndicBERTroBERTaHingBERTBERTBanglishBERTmBERTMuRILXLM-RTri-Distil-BERTMixed-Distil-BERT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.590.590.590.590.60.60.60.60.660.660.660.660.660.660.660.660.660.660.660.660.670.670.670.670.680.680.680.680.720.720.720.720.740.740.740.740.770.770.770.770.770.770.770.770.690.690.690.690.70.70.70.7Weighted Weighted F1 Score
@@
HindiBERTBanglaBERTDistilBERTroBERTafBERTHateBERTMuRILIndicBERTHingBERTBERTBanglishBERTmBERTXLM-RTri-Distil-BERTMixed-Distil-BERT000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.730.730.730.730.760.760.760.760.790.790.790.790.80.80.80.80.810.810.810.810.810.810.810.810.810.810.810.810.820.820.820.820.820.820.820.820.830.830.830.830.860.860.860.860.880.880.880.880.880.880.880.880.860.860.860.860.870.870.870.87Weighted Weighted F1 Score